# Human Attention During Localization of Memory Bugs in C Programs

### v1

@@
Abstract
This paper presents a study of human visual attention during localization of memory bugs in C. Human visual attention refers to the mechanical processes by which we selectively process and prioritize information. Visual attention is important to study because it is central to what information people (who are sighted) use to solve a particular problem. Meanwhile, memory bugs are among the most common types of bugs in C programs that manifest as a variety of program faults. In this paper, we study human visual attention while people attempt to locate memory bugs in code. We recruit 11 programmers to locate between one and six memory bugs in three C programs for 1.5-2 hours each. In total we collected observations of 17 hours of programmer effort. The bugs in our study cover memory leaks, overflows, and double frees, which are among the most common memory bugs. We analyze the task outcomes in terms of success rate and related factors, patterns of visual attention overall such as what lines and functions are read, and finally we explore differences of visual attention patterns during success versus failure cases.
Index Terms:
@@ Abstract
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
1 Introduction
@@
This paper presents a study of human attention during localization of C memory bugs. Human visual attention refers to the mechanical processes by which we selectively process and prioritize information [1]. Human attention can be measured through eye tracking, interface logs, and activity surveys. It is important to study because beliefs about human visual attention influence decisions such as personnel allocation, user interface design, and even the architecture of machine neural networks [2, 3, 4, 5, 6, 7, 8]. Human attention has also long been academically interesting for “greater” philosophical reasons as a window into the working of our own minds [9].
@@
Meanwhile, memory bugs are one of the most common programming errors in C programs [10, 11]. Memory bugs are behind aberrant software behavior such as excessive memory use, slowdowns, security vulnerabilities, and crashes. Some memory bug behaviors such as buffer overflows and segmentation faults are so common as to be tropes in C programming lore [12, 13]. Yet while often the cause of a memory bug is relatively small—a missing free(), an array length miscalculation, etc.—memory bugs can be difficult for programmers to find. Complex programs can have many pointers active at a time and memory leaks can be hard to observe as they accumulate slowly in long-running programs. This difficulty has made memory bugs a risk in C programs and a major research focus [10, 14].
@@
Three types of memory bug dominate. Cotroneo et al. [15] point out that “use of uninitialized data, buffer overflows, and memory leaks are preponderant” as causes of bug behavior related to memory. Part of what makes these difficult to find is that they can lead to intermittent problems or challenges with vague non-functional requirements like responsiveness. For example it may be possible to use uninitialized data for a long time before it causes a problem, even if it is not strictly legal. Memory leaks may manifest as slow degradation in user experience rather than a diagnosable error message. Better understanding how people find these types of bugs could help address the lion’s share of challenges in localizing C memory bugs.
@@
In this paper, we conduct an experiment in which programmers read C code and find memory bugs. We extracted eight bug reports from three large, popular C programs. Each bug is related to the three dominant memory bug types described in the previous paragraph or was a common bug type in the open-source C repositories. We recruited 11 programmers; in the study each programmer worked for 1.5 to 2 hours to find as many of the six presented bugs as possible. During the study we used eye tracking hardware to observe each programmer’s visual attention, while also tracking other interactions. We measured programmer performance such as success rate in finding bugs and time required. We also used a minimally-invasive attentional state survey for a self-assessment of attentiveness. Taken together these data form a view of visual attention during the bug localization tasks in C programs.
@@
2 Background and Related Work
@@
2.1 Studies of Human Attention
@@
Eye tracking has been used extensively to study human attention. For example, computational models of eye movement during reading have been developed to predict how a person’s eyes will move over a body of text [16, 17]. These models highlight that eye movement across text can be predicted by textual features, which cause people to move from word to word, but also pause and reread text to ensure understanding. Such models have also been applied to non-reading tasks, such as visual search [18]. Collectively, the modeling efforts have revealed that there are systematic patterns of eye movements that people follow when reading or performing other visual tasks. Consequently, eye tracking has also been employed to examine when and why people deviate from those stereotypical patterns, thereby identifying the oculomotor correlates of mindlessness (or mind-wandering) during reading and other tasks. Specifically, multiple studies have found systematic differences in eye movements when people are reading attentively vs. inattentively [19, 20, 21, 22]. Specifically, when people are mind-wandering while reading, their eye movements are less sensitive to the content and complexity of the text, moving over the text unpredictably and unsystematically.
@@
These phenomena have also been studied in other contexts, such as watching video lectures to simulate online coursework [23], listening to audiobooks [24], and viewing images of scenes [24, 25]. Faber et al. [24] have highlighted that inattentiveness is not always accompanied by the same patterns of eye movements across various tasks, as they require differences in the focality/dispersion of visual attention. Therefore, the task itself must be considered when examining such differences. The present study will thus seek to identify systematic oculomotor patterns when people are localizing bugs in code in an effort to understand how people do and do not detect bugs, and what their eyes are doing when that happens.
2.2 Eye tracking in Program Comprehension
@@
Eye tracking technology has been a bedrock of program comprehension research for at least two decades and recent systematic surveys have led to recommended best practices in using this technology [26, 27]. Over time simultaneous hardware cost decreases and quality increases have expanded access to eye tracking equipment and the number of papers using it [28].
@@
Program comprehension itself has a long history as research discovering how people read and understand source code [29, 30, 31, 32, 33]. Sometimes formalized as the “concept assignment problem” [34], the idea is that reading code is the process of interpreting the low-level implementation details in code as high-level concepts about program behavior. Eye tracking helps investigate the mechanical process of reading those low-level details. For example, Aschwanden and Crosby [35] observed how programmers’ eyes moved while examining a snippet of code featuring a loop and equations. Bednarik and Tukiainen [36] focused on scanpaths (i.e., sequences of eye movements) as participants tried to make sense of two small programs. Their findings revealed that programmers’ mental strategies sharpen over time, as indicated by fewer shifts in visual focus. Later, Rodeghero et al. [37] turned their attention to how developers write documentation, using this information to inform advances in automated summarization tools. Followup examination of these results extracted and categorized different eye tracking movement strategies [38, 39]. Abid et al. [40] expanded on Rodeghero’s study [37], incorporating real-world project contexts and a realistic IDE interface (better interfaces are possible with newer eye tracking devices and infrastructure). Similar findings have been observed in related areas such as reading UML diagrams, with some diagram formats leading to more thorough program comprehension than others [41, 42].
@@
A few papers have sought to combine programmer visual attention with neural network-based models to predict human behavior or change machine behavior to be more like human behavior. For example, Karas et al. [43] compared eye movement pattern strategies when reading code for different tasks, then used these results to design novel neural network architectures to do those tasks [5]. Likewise Bansal et al. [44] modeled eye movements as scanpaths and augmented the attention mechanism in a neural network to improve code documentation generation [45]. The idea of combining human with AI attention is starting to take root outside of software engineering, with preliminary evidence showing performance improvements in various tasks such as image recognition and text generation [46, 47].
2.3 Studies of Bug Localization
@@
Bug localization is the task of finding the area of code that causes a bug behavior [48, 49, 50]. Bug localization is a popular research target because understanding how people find bugs could lead to better automated tools to help them. Studies of bug localization have revealed strategies roughly falling into either static or dynamic approaches. Static strategies are those reading code itself while dynamic approaches execute that code. Dynamic approaches are well-studied and include tactics like program logging, assertions, breakpoints, and profiling [51]. Static approaches are less well-studied, though are a common starting point for programmers, especially when the code exists in a large codebase with which the programmer is not yet familiar. This situation is common when onboarding new staff, for example [52, 53]. This paper is primarily concerned with static approaches and specifically the mechanical process of attention during bug localization. Related papers include eye tracking studies such as by Bednarik and Tukiainen [54] who analyze the effect of experience level, finding that people with more experience read more areas of code but in less depth than novices. Research indicates that eye movements differ between debugging code and reading for other tasks [55]. One difference is that programmers search code more often in some tasks than others [56]. Given that differences exist in different tasks, this paper focuses on the narrow task of localization of memory bugs in C programs to avoid an assumption that movement patterns for general tasks generalize to this specific task.
3 Experimental Design
@@ 3.1 Research Questions
3.2 Procedure Overview
Our experimental procedure is a single-group cohort [57], “Quadrant II” study described by Stol and Fitzgerald [58]. What this means is that we focus on a single-group common characteristic (people proficient in C programming) in a controlled environment (as opposed to “Quadrant I” field studies, for example). We chose a single-group design because we do not seek to distinguish human factors such as age or gender, and we chose a controlled environment to reduce experimental variables. We 1) curate a list of bugs that is representative of typical memory bugs in C programs, 2) recruit programmers proficient in C to read bug reports and source code from that list, and 3) observe what the people do in a controlled environment.
@@ 3.2 Procedure Overview
We curated a pool of eight bugs and we recruited 11 people. We present each person a maximum of six bugs to find over 1.5 to 2 hours. Three bugs were the same for all participants and three rotated from our pool. To reduce experimental variables, we conducted the study one person at a time, with each person physically present in a quiet computer laboratory setting using the same hardware and software. Note that because we are primarily interested in patterns of attention, we avoided distracting participants from the computer. We chose a minimally-invasive design where an author of this paper was available for experimental setup at all times, but this person did not intervene or ask the participant questions in an e.g. think-aloud form.
3.3 Subject Bug Reports
We selected a total of eight bug reports from three C libraries/programs. Our goal is for these bug reports to be representative of typical memory bugs in C. For a definition of “typical” we turned to Cotroneo et al. [15] who studied characteristics of bugs and found that memory bugs are almost always one of three types: uninitialized data, buffer overflows, and memory leaks. We use bugs of the buffer overflow and memory leak type, but used double frees in lieu of uninitialized data due to their prevalence in our datasets. Bugs also have a range of difficulty and severity. At the same time we must balance diversity of bug selection with the practical reality of human studies, as has been pointed out in retrospective studies e.g. by Siegmund et al. [59]. Participants will learn about a program and seek to reuse that knowledge on other bugs, they will have different prior knowledge, have time limits, and will fatigue.
Table I below shows the three C libraries/programs we used. These programs have a range of total size and domain. Total size is important because finding bugs in larger programs may be more difficult. Domain is important because coding style may differ depending on domain [60]. We then selected between one and three bugs from the bug report repository for each program. We selected easy, medium, and hard difficulty bugs, where we assessed difficulty using comments in the bug report conversation, area of code that needed to be changed to fix the bug, and our own judgment. We also ensured that all bugs were double-frees, buffer overflows, or memory leaks. See Section 3.7 for a discussion of the risks and limitations of our choice of bugs. Also see our reproducibility package linked in Section 5.
@@ 3.3 Subject Bug Reports
TABLE I: Statistics for C libraries from which we selected bugs.
3.4 Study Participants
@@
We recruited 11 participants for this study. The participants had between 3.5 and 10 years experience programming. Five were graduate students and six were undergraduates. The undergraduates were recommended as strong C programmers by faculty, though these faculty were not authors of this study and this study was not connected to any coursework (per IRB requirements). We compensated participants at the fully-loaded local market rate of US\$100/hr.
3.5 Procedure Details
@@
The details of our study procedure boil down to: 1) each participant visited the laboratory in person at a scheduled time, one at a time, 2) the participant sat at the eye tracking computer to become familiar with the settings and review informed consent documents, and 3) read and locate C memory bugs for a maximum of 120 minutes. We gave each participant a maximum of six bug reports to read. Some of the bug reports were the same for everyone, while we rotated others (as a balance between diversity of samples and measuring consistency of participant behavior). The first two bug reports were the same for all participants (see online appendix Section 5.1.2). The third, fourth, fifth, and sixth bug reports rotated from the pool of eight.
@@
Figure 1: The interface presented to participants. Area A is the bug report, area B is the code, area C is the file explorer, and area D is a text chat window with GPT-4o for AI help.
@@
An example image of the interface participants saw is in Figure 1. To reduce experimental variables and ensure accurate eye tracking, we did not allow the participants to navigate away from the interface; we instead made necessary information available within the interface. The interface itself is based on the Eclipse 4.33.0 IDE with iTrace 0.2.0 [61]. We set the IDE to have four areas, shown in Figure 1. Area A is information about the bug including the bug report. We showed only the portion of the report explaining the bug symptoms. We did not allow the participants to see followup in which the exact bug location may be revealed. Area A is also where participants wrote their answers.
@@
We regulated access to outside resources via a chat window with GPT-4o using the RemainAI [62] Eclipse plugin (Area D in Figure 1). This window was necessary because: 1) we did not allow focus to be changed from the IDE interface, 2) we did not want participants to find the bug’s solution from a solved bug report online, 3) in practice AI help is quickly becoming an important part of the programming process so participants expect this support, and 4) the plugin allows us to capture queries that people type. We set the plugin so that it would not automatically have access to the code and we asked participants to ask general coding help questions instead of details about the program. This restriction was necessary because GPT-4o could have memorized the bug report information during training.
@@
The purpose of this window is to measure self-reported attention level. This method is used in cognitive psychology to generate momentary subjective assessments of attentional states [63, 64, 65]. Responses to these screens can be used to examine which signals, both in terms of overt behavioral responses and physiological indices, co-occur with and precede moments of attentiveness vs. inattentiveness. Similarly, in the present study, we sought to understand whether certain behaviors/eye tracking metrics would be associated with self-reported inattentiveness by participants.

### v2

@@
Abstract.
This paper presents a study of human visual attention during localization of memory bugs in C. Human visual attention refers to the mechanical processes by which we selectively process and prioritize information. Visual attention is important to study because it is central to what information people (who are sighted) use to solve a particular problem. Meanwhile, memory bugs are among the most common types of bugs in C programs that manifest as a variety of program faults. In this paper, we study human visual attention while people attempt to locate memory bugs in code. We recruit 21 programmers to locate between one and eight memory bugs in three C programs for 1.5–2 hours each. In total we collected observations of 31 hours of programmer effort. The bugs in our study cover memory leaks, overflows, and double frees, which are among the most common memory bugs. We analyze the task outcomes in terms of success rate and related factors, patterns of visual attention overall such as what lines and functions are read, and finally we explore differences of visual attention patterns during success versus failure cases.
@@ Abstract.
copyright: acmlicensedjournal: TOSEMjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Software development techniquesccs: Social and professional topics Computing educationccs: Human-centered computing Empirical studies in HCI
1. Introduction
@@
This paper presents a study of human attention during localization of C memory bugs. Human visual attention refers to the mechanical processes by which we selectively process and prioritize information (Carrasco, 2011). Human attention can be measured through eye tracking, interface logs, and activity surveys. It is important to study because beliefs about human visual attention influence decisions such as personnel allocation, user interface design, and even the architecture of machine neural networks (Novák et al., 2024; Baharum et al., 2024; Cheng and Liu, 2012; Zhang et al., 2024, 2025; Pourhosein and Sabokro, 2025; Kiseleva et al., 2020). Human attention has also long been academically interesting for “greater” philosophical reasons as a window into the working of our own minds (Mole, 2025).
@@
Meanwhile, memory bugs are one of the most common programming errors in C programs (Van der Veen et al., 2012; The Chromium Project, [n. d.]). Memory bugs are behind aberrant software behavior such as excessive memory use, slowdowns, security vulnerabilities, and crashes. Some memory bug behaviors such as buffer overflows and segmentation faults are so common as to be tropes in C programming lore (Durumeric et al., 2014; Orman, 2003). Yet while often the cause of a memory bug is relatively small—a missing free(), an array length miscalculation, etc.—memory bugs can be difficult for programmers to find. Complex programs can have many pointers active at a time and memory leaks can be hard to observe as they accumulate slowly in long-running programs. This difficulty has made memory bugs a risk in C programs and a major research focus (Van der Veen et al., 2012; van Oorschot, 2023).
@@
Cotroneo et al. (Cotroneo et al., 2016) point out that “use of uninitialized data, buffer overflows, and memory leaks are preponderant” as causes of bug behavior related to memory. Part of what makes these difficult to find is that they can lead to intermittent problems or challenges with vague non-functional requirements like responsiveness. For example it may be possible to use uninitialized data for a long time before it causes a problem, even if it is not strictly legal. Memory leaks may manifest as slow degradation in user experience rather than a diagnosable error message. Better understanding how people find these types of bugs could help address the lion’s share of challenges in localizing C memory bugs.
@@
In this paper, we conduct an experiment in which programmers read C code and find memory bugs. We extracted eight bug reports from three large, popular C programs. Each bug is related to the three dominant memory bug types described in the previous paragraph or was a common bug type in the open-source C repositories. We recruited 21 programmers; in the study each programmer worked for 1.5 to 2 hours to find as many of the eight presented bugs as possible. During the study we used eye tracking hardware to observe each programmer’s visual attention, while also tracking other interactions. We measured programmer performance such as success rate in finding bugs and time required. We also used a minimally-invasive attentional state survey for a self-assessment of attentiveness. Taken together these data form a view of visual attention during the bug localization tasks in C programs.
@@
2. Background and Related Work
@@
2.1. Studies of Human Attention
@@
Eye tracking has been used extensively to study human attention. For example, computational models of eye movement during reading have been developed to predict how a person’s eyes will move over a body of text (Reichle et al., 2003; Engbert et al., 2005). These models highlight that eye movement across text can be predicted by textual features, which cause people to move from word to word, but also pause and reread text to ensure understanding. Such models have also been applied to non-reading tasks, such as visual search (Reichle et al., 2012). Collectively, the modeling efforts have revealed that there are systematic patterns of eye movements that people follow when reading or performing other visual tasks. Consequently, eye tracking has also been employed to examine when and why people deviate from those stereotypical patterns, thereby identifying the oculomotor correlates of mindlessness (or mind-wandering) during reading and other tasks. Specifically, multiple studies have found systematic differences in eye movements when people are reading attentively versus inattentively (Reichle et al., 2010; Foulsham et al., 2013; Frank et al., 2015; Bixler and D’Mello, 2016). Specifically, when people are mind-wandering while reading, their eye movements are less sensitive to the content and complexity of the text, moving over the text unpredictably and unsystematically.
@@
These phenomena have also been studied in other contexts, such as watching video lectures to simulate online coursework (Zhang et al., 2020), listening to audiobooks (Faber et al., 2020), and viewing images of scenes (Faber et al., 2020; Krasich et al., 2018). Faber  et al(Faber et al., 2020) have highlighted that inattentiveness is not always accompanied by the same patterns of eye movements across various tasks, as they require differences in the focality/dispersion of visual attention. Therefore, the task itself must be considered when examining such differences. The present study will thus seek to identify systematic oculomotor patterns when people are localizing bugs in code in an effort to understand how people do and do not detect bugs, and what their eyes are doing when that happens.
2.2. Eye tracking in Program Comprehension
@@
Eye tracking technology has been a bedrock of program comprehension research for at least two decades and recent systematic surveys have led to recommended best practices in using this technology (Sharafi et al., 2015, 2020b). Over time simultaneous hardware cost decreases and quality increases have expanded access to eye tracking equipment and the number of papers using it (Braw et al., 2023).
@@
Program comprehension itself has a long history as research discovering how people read and understand source code (Cornelissen et al., 2009; Dominic et al., 2020b; Maalej et al., 2014; Schröter et al., 2017; Stapleton et al., 2020). Sometimes formalized as the “concept assignment problem” (Biggerstaff et al., 1993), the idea is that reading code is the process of interpreting the low-level implementation details in code as high-level concepts about program behavior. To study how developers engage with these low-level details, researchers often use eye tracking to examine the mechanics of visual attention. One fundamental component of this analysis is fixations, when the eyes briefly pause on a specific point to allow the brain to process visual information (Grabinger et al., 2025). Analyzing fixations provides insight into how developers allocate attention during code reading. For example, Aschwanden and Crosby (Aschwanden and Crosby, 2006) observed how programmers’ eyes moved while examining a snippet of code featuring a loop and equations. Bednarik and Tukiainen (Bednarik and Tukiainen, 2006) focused on scanpaths (i.e., sequences of eye movements) as participants tried to make sense of two small programs. Their findings revealed that programmers’ mental strategies sharpen over time, as indicated by fewer shifts in visual focus. Peitek  et al. showed that knowledge guides the eyes of proficient programmers (Peitek et al., 2022). Later, Rodeghero et al. (Rodeghero et al., 2014) turned their attention to how developers write documentation, using this information to inform advances in automated summarization tools. Followup examination of these results extracted and categorized different eye tracking movement strategies (Rodeghero and McMillan, 2015; Rodeghero et al., 2015). Abid et al. (Abid et al., 2019) expanded on Rodeghero’s study (Rodeghero et al., 2014), incorporating real-world project contexts and a realistic IDE interface (better interfaces are possible with newer eye tracking devices and infrastructure). Similar findings have been observed in related areas such as reading UML diagrams, with some diagram formats leading to more thorough program comprehension than others (Guéhéneuc, 2006; Sharif and Maletic, 2010).
@@
A few papers have sought to combine programmer visual attention with neural network-based models to predict human behavior or change machine behavior to be more like human behavior. For example, Karas et al. (Karas et al., 2024) compared eye movement pattern strategies when reading code for different tasks, then used these results to design novel neural network architectures to do those tasks (Zhang et al., 2024). Likewise Bansal et al. (Bansal et al., 2023b) modeled eye movements as scanpaths and augmented the attention mechanism in a neural network to improve code documentation generation (Bansal et al., 2023a). The idea of combining human with AI attention is starting to take root outside of software engineering, with preliminary evidence showing performance improvements in various tasks such as image recognition and text generation (Fong et al., 2018; Lopez-Cardona et al., 2024).
2.3. Studies of Bug Localization
@@
Bug localization is the task of finding the area of code that causes a bug behavior (Saha et al., 2013; Thung et al., 2014; Obaidellah et al., 2018). Bug localization is a popular research target because understanding how people find bugs could lead to better automated tools to assist people. Studies of bug localization have revealed strategies roughly falling into either static or dynamic approaches. Static strategies are those reading code itself while dynamic approaches execute that code. Dynamic approaches are well-studied and include tactics like program logging, assertions, breakpoints, and profiling (Wong et al., 2016). Static approaches are less well-studied, though are a common starting point for programmers, especially when the code exists in a large codebase with which the programmer is not yet familiar. This situation is common when onboarding new staff, for example (Dominic et al., 2020a; Ju et al., 2021). This paper is primarily concerned with static approaches and specifically the mechanical process of attention during bug localization. Related papers include eye tracking studies such as by Bednarik and Tukiainen (Bednarik and Tukiainen, 2008) who analyze the effect of experience level, finding that people with more experience read more areas of code but in less depth than novices. Similarly to results from program comprehension studies, Hejmady and Narayanan analyzed scanpaths during debugging tasks and found that better programmers have fewer attentional shifts (Hejmady and Narayanan, 2012). Research indicates that eye movements differ between debugging code and reading for other tasks (Turner et al., 2014). One difference is that programmers search code more often in some tasks than others (Sharafi et al., 2020a). Given that differences exist in different tasks, this paper focuses on the narrow task of localization of memory bugs in C programs to avoid an assumption that movement patterns for general tasks generalize to this specific task.
3. Experimental Design
@@ 3.1. Research Questions
3.2. Procedure Overview
Our experimental procedure is a single-group cohort (Kazdin, 2016), “Quadrant II” study described by Stol and Fitzgerald (Stol and Fitzgerald, 2020). What this means is that we focus on a single-group common characteristic (people proficient in C programming) in a controlled environment (as opposed to “Quadrant I” field studies, for example). We chose a single-group design because we do not seek to distinguish human factors such as age or gender, and we chose a controlled environment to reduce experimental variables. We 1) curate a list of bugs that is representative of typical memory bugs in C programs, 2) recruit programmers proficient in C to read bug reports and source code from that list, and 3) observe what the people do in a controlled environment.
@@ 3.2. Procedure Overview
We curated a pool of eight bugs and we recruited 21 people. We present each person a maximum of eight bugs to find over 1.5 to 2 hours. Three bugs were the same for all participants and three to five rotated from our pool. To reduce experimental variables, we conducted the study one person at a time, with each person physically present in a quiet computer laboratory setting using the same hardware and software. Note that because we are primarily interested in patterns of attention, we avoided distracting participants from the computer. We chose a minimally-invasive design where an author of this paper was available for experimental setup at all times, but this person did not intervene or ask the participant questions in an e.g. think-aloud form. The scope of this paper is a quantitative analysis of the eye tracking data.
Table 1. Statistics for C libraries from which we selected bugs.
@@
3.3. Subject Bug Reports
We selected a total of eight bug reports from three C libraries/programs. Our goal is for these bug reports to be representative of typical memory bugs in C. For a definition of “typical” we turned to Cotroneo et al. (Cotroneo et al., 2016) who studied characteristics of bugs and found that memory bugs are almost always one of three types: uninitialized data, buffer overflows, and memory leaks. We use bugs of the buffer overflow and memory leak type, but used double frees in lieu of uninitialized data due to their prevalence in our datasets. Bugs also have a range of difficulty and severity. At the same time we must balance diversity of bug selection with the practical reality of human studies, as has been pointed out in retrospective studies e.g. by Siegmund et al. (Siegmund et al., 2015). Participants will learn about a program and seek to reuse that knowledge on other bugs, they will have different prior knowledge, have time limits, and will fatigue.
Table 1 above shows the three C libraries/programs we used. These programs have a range of total size and domain. Total size is important because finding bugs in larger programs may be more difficult. Domain is important because coding style may differ depending on domain (Cox and Fisher, 2009). We then selected between one and three bugs from the bug report repository for each program. We selected easy, medium, and hard difficulty bugs, where we assessed difficulty using comments in the bug report conversation, area of code that needed to be changed to fix the bug, and our own judgment. We also ensured that all bugs were double-frees, buffer overflows, or memory leaks. See Section 3.7 for a discussion of the risks and limitations of our choice of bugs. Also see our reproducibility package linked in Section 5.
Figure 1. This is the bug report for ladybug. This report is adapted from issue #6861 in the sway repository.:
3.4. Study Participants
@@
We recruited 21 participants for this study. The participants had between 3 and 13 years of programming experience. Eleven were graduate students, nine were undergraduates, and one was a professional software engineer. Eight of the graduate students worked in the software engineering industry before coming to graduate school. Seven of the undergraduates had professional experience at software engineering internships. The undergraduates were recommended as strong C programmers by faculty, though these faculty were not authors of this study and this study was not connected to any coursework (per IRB requirements). We compensated participants at the fully-loaded local market rate of 100 USD/hr.
3.5. Procedure Details
@@
The details of our study procedure boil down to: 1) each participant visited the laboratory in person at a scheduled time, one at a time, 2) the participant sat at the eye tracking computer to become familiar with the settings and review informed consent documents, and 3) read and locate C memory bugs for a maximum of 120 minutes. We gave each participant a maximum of eight bug reports to read. Some of the bug reports were the same for everyone, while we rotated others (as a balance between diversity of samples and measuring consistency of participant behavior). The first two bug reports were the same for all participants (see online appendix Section 5.1.2). The remaining reports rotated from the pool of eight.
@@
Figure 2. The interface presented to participants. Area A is the bug report, area B is the code, area C is the file explorer, and area D is a text chat window with GPT-4o for AI help.:
@@
An example image of the interface participants saw is in Figure 2. To reduce experimental variables and ensure accurate eye tracking, we did not allow the participants to navigate away from the interface; we instead made necessary information available within the interface. The interface itself is based on the Eclipse 4.33.0 IDE with iTrace 0.2.0 (Zyrianov et al., 2020). We set the IDE to have four areas, shown in Figure 2. Area A is information about the bug including the bug report. We showed only the portion of the report explaining the bug symptoms. We did not allow the participants to see followup in which the exact bug location may be revealed. Area A is also where participants wrote their answers.
@@
We regulated access to outside resources via a chat window with GPT-4o using the RemainAI (Remain Software, 2025) Eclipse plugin (Area D in Figure 2). This window was necessary because: 1) we did not allow focus to be changed from the IDE interface, 2) we did not want participants to find the bug’s solution from a solved bug report online, 3) in practice AI help is quickly becoming an important part of the programming process so participants expect this support, and 4) the plugin allows us to capture queries that people type. We set the plugin so that it would not automatically have access to the code and we asked participants to ask general coding help questions instead of details about the program. This restriction was necessary because GPT-4o could have memorized the bug report information during training.
@@
The purpose of this window is to measure self-reported attention level. This method is used in cognitive psychology to generate momentary subjective assessments of attentional states (Weinstein, 2018; Robison et al., 2019; Smallwood and Schooler, 2015). Responses to these screens can be used to examine which signals, both in terms of overt behavioral responses and physiological indices, co-occur with and precede moments of attentiveness versus inattentiveness. Similarly, in the present study, we sought to understand whether certain behaviors/eye tracking metrics would be associated with self-reported inattentiveness by participants.